OpenStudy (anonymous):

Determine which of the following formulas gives the nth term of this sequence. 38, 40 ,42, 44, 46,....... 36 + 2n 40 - 2n 39 - n 37 + n

OpenStudy (anonymous):

All of the answers give 38 when n=1....as they should Which answer gives 40 when n =2 ?

OpenStudy (anonymous):

huuuuuh

OpenStudy (anonymous):

right here

OpenStudy (anonymous):

38 40 42 44 46

OpenStudy (anonymous):

this is a arithmatic progression

OpenStudy (anonymous):

first term=a=38 common difference= 40-38=2 nth term=a+(n-1)*d =38+(n-1)2 =38+2n-2 =2n+36

OpenStudy (anonymous):

these are the answer choices